# Sensitivity and topology of exceptional rings in nonlinear non-Hermitian planar optical microcavities

### v1

@@ III Nonlinearity-induced third-order exceptional ring
We begin the discussion of the nonlinear eigenvalue problem for the case of zero detuning, $\Delta=0$, showing eigenvalues and further analysis in Figs. 3 and 4. As shown in Appendix B, we have two solutions in region A and C, and four solutions in region B. In the following figures we show, for clarity, eigenvalues $\mu^{\prime}=\mu-\mu_{0}$, with $\mu_{0}$ defined in Appendix B. Figure 3 shows the resulting real and imaginary parts of the complex-valued energy spectrum and central cuts for $\alpha=0.25~{}\mathrm{meV\upmu m}^{2}$. In the nonlinear regime the complex-valued energy spectrum hosts two concentric rings marked by solid blue and violet dots in the central cuts in Figure 3(c-d) separating three radial segments (A-C). The inner segment $A$ hosts two circularly polarized solutions, the segment $B$ between the rings, marked in red, features four solutions (two circularly and two linearly polarized), and the outer segment $C$ shows two linearly polarized solutions. In Figure 4(a-b) these segments are sketched for an arbitrary nonlinearity $\alpha\neq 0$ to visualize the bifurcation behavior at the two rings. Analyzing this behavior, we conclude that in the nonlinear regime the linear ER splits into an inner ring of fold points (blue dot) along $|\textbf{k}|^{2}=\nicefrac{{|\gamma|}}{{|\Delta_{\mathrm{LT}}|}}$ and an outer ring of cusp points (violet dot) along $\sqrt{\nicefrac{{(\alpha^{2}+\gamma^{2})}}{{\Delta_{\mathrm{LT}}^{2}}}}$, which grows with increasing Kerr-like nonlinearity $g_{\mathrm{c}}$ or cross-interaction strength $g_{\mathrm{x}}$. The fold and cusp points can be identified with exceptional points of $2^{\mathrm{nd}}$ and $3^{\mathrm{rd}}$-order, respectively. Hence, the outer ring is a $3^{\mathrm{rd}}$-order exceptional ring. The circles mark points where a solution crosses the boundary between two segments without tracing through an ER (e.g. the two green colored solutions form an ER at the blue solid dot, at the border of segment A and B, while the blue and red colored solutions cross the border of segment A and B at the blue circle without forming an ER). For $\alpha<0$ the cusp ring is redshifted in relation to the fold ring [this scenario is displayed in Figure 1(a)]. While previous work realized individual higher-order exceptional points in two-dimensional nonlinear parameter space and thus reduced the parameter tuning requirements [71, 70], the ring structure observed here reduces these parameter tuning requirements even further, making higher-order exceptional point structures in low dimensions more accessible.
At this point, we need to discuss the nature of exceptional points of order $n>2$ in nonlinear $2\times 2$ systems. In previous work it was exemplarily shown that mapping a nonlinear and non-Hermitian Hamiltonian to a $n\times n$ linear and $\mathcal{PT}$-symmetric Hamiltonian features the same exceptional point of order $n$ [71, 70]. Accordingly, the singularity in the nonlinear system is called a “nonlinear exceptional point”, containing at least one dynamically unstable solution. It is straightforward to show that the four solutions in section B in Figure 3(c-d) form an overcomplete basis. This is reflected in the Petermann factor $\mathrm{PF}=\left|\braket{\psi_{L}}{\psi_{R}}\right|^{-2}$, which serves as a measure for non-orthognality of left $\bra{\psi_{L}}$ | and right |$\ket{\psi_{R}}$ eigenvectors [82]. $\mathrm{PF}=1$ indicates orthogonal modes, whereas $\mathrm{PF}>1$ signals mode non-orthogonality, resulting in a divergent Petermann factor at linear exceptional points. At low nonlinearity the nonlinear $3^{\mathrm{rd}}$-order ER exhibits diverging $\mathrm{PF}$. With increasing nonlinearity the Petermann factor at the nonlinear $3^{\mathrm{rd}}$-order ER decreases until it converges to $1$. This means that the eigenvectors at the ER become increasingly orthogonal and do not coincide in the nonlinear regime. However, it has been shown that nonlinear exceptional points still feature an intriguing spectral topology and the $n^{\mathrm{th}}$-root perturbation response of their linear counterparts [23]. This suggests that the unstable mode still contributes to the order of the nonlinear exceptional point [71, 70]. In that sense we believe that the nonlinear ER appears to show sufficient similarities with its linear counterpart to justify its name and order classification.
@@
Figure 4: Elliptic umbilic structure of exceptional rings and exceptional points in nonlinear regime. (a) Sketch of the three sections A-C [as displayed in Fig. 3(c)] with two, four, and two solutions, respectively. (b) The bifurcation behavior on the edges of ring B characterizes the inner (outer) ring as a ring of fold (cusp) points. (c) Radii of fold rings (turquoise lines) and cusp-rings (violet dots) for degnerate modes (segments marked as above) and for finite mode detuning $\Delta$. The umbilic structure shows, that fold-rings are part of exceptional arcs and cusp-rings are located at exceptional nexuses. (d) Cross section in (c) for variable nonlinearity $\alpha$. The depicted elliptic umbilic predicts exceptional ring radii and type in the full parameter space. The central organization point represents a point on the linear exceptional ring discussed in Section II.
@@
Figure 5: Enhanced perturbation response at nonlinear exceptional rings and points. In (a-b) a $3^{\mathrm{rd}}$-order exceptional point on the exceptional nexus ring [star in (d)] is traced for increasing $\alpha$. (a) The inverse signal-enhancement factor $(\mathrm{SEF})^{-1}$ decreases in the nonlinear regime. The coloring visualizes $\alpha$ in each curve (growing from black to yellow). (b) The maximum $\mathrm{SEF}_{\mathrm{max}}$ at the exceptional point as a function of $\alpha$. (c) Response $\Delta_{\omega}$ of the exceptional points on the exceptional arc ring (EAR) and exceptional nexus ring (EXR) on the elliptic umbilic along $\Delta=0$ [horizontal dashed line in (d) corresponds to $\epsilon=0$ in (c)] for a fixed nonlinearity $\alpha$. The response of the EAR significantly deviates from the square-root law of linear exceptional points. Here, small perturbations tune the system along the left exceptional arc of the umbilic [red line in (d)]. Larger perturbations tune it away from the arc due to its curvature [blue area in (d)] leading to growing response. (d) Colored cut through elliptic umbilic at finite nonlinearity $\alpha$. The coloring denotes the response function scaling factor $\eta$ when perturbing an exceptional point on the umbilic singularity structure. Detuning allows to realize different rings on the umbilic, which feature distinct perturbation responses. Perturbation from the lower to the upper arc (or vice versa) leads to characteristic square-root response $\eta_{1}$. Nexus perurbation leads to stronger response $\eta_{2}$. The perturbation $\epsilon$ is defined in relation to the respective point on the umbilic.
In previous work [37] it was mentioned that the linear ER does not exist in the solutions of the eigenvectors for finite mode detuning. Here we have shown that in the nonlinear regime new ERs arise even for finite mode detuning. We use this knowledge to analyze the perturbation response of nonlinear ERs in Section IV and their topological features under mode detuning in Section V. In Appendix C we detail that the elliptic umbilic structure applies to a wide range of non-Hermitian Hamiltonians when introducing nonlinearity. Therefore, the topolgoical features and perturbation response of nonlinear ERs on the umbilic we discuss below, are generic and can be applied to exceptional points in many other systems and constellations.
@@ IV Enhanced Perturbation Response at Nonlinear Exceptional Rings and Points
In Ref. [71, 88] it was proposed that nonlinear exceptional points can exhibit a complete basis of eigenstates. In laser systems, the Petermann factor can be investigated as a measure of noise enhancement. Based on this, it was proposed in Ref. [70] that nonlinear exceptional points can exhibit a divergent signal-to-noise ratio. More recently, however, it was shown that the conventional Petermann factor does not cover the interplay of noise and nonlinearity and that nonlinear exceptional points exhibit finite signal-to-noise ratio [88], with a generalized form of the Petermann factor at exceptional points derived in Ref. [89]. In the following, we discuss that nonlinear exceptional points possess adjustable and in some cases ultra-enhanced perturbation response. Our analysis covers the group of both real- and imaginary-valued smooth nonlinearities that lead to the elliptic umbilic exceptional point structure described above; here exemplarily evaluted for Kerr-type coherent nonlinearities. The elliptic umbilic structure significantly increases the complexity of an exceptional points neighborhood permiting sensor design for exceptional points at different positions on the singularity structure, which changes the response to perturbations not limited to $n^{\mathrm{th}}$-root laws.
@@
In Figure 5(b) the maximum of the signal-enhancement factor $\mathrm{SEF}_{\mathrm{max}}$ is plotted against $\alpha$. The maximum signal enhancement of $\mathrm{SEF}=1.9\times 10^{9}$ is found for a nonlinearity strength ${(g_{\mathrm{c}}-g_{\mathrm{x}})}/{2}=68~{}\mathrm{\upmu eV\upmu m^{2}}$. Here, the signal enhancement scales with the perturbation as $\mathrm{SEF}\propto|\epsilon|^{-1.29}$. In contrast to an isolated linear exceptional point or ER, the elliptic umbilic allows for different exceptional points to be perturbed depending on the initial detuning $\Delta$. Since the perturbation $\epsilon$ is parallel to $\Delta$ this means that depending on which exceptional point we chose to perturb, the response depends on its position on the umbilic. In planar nonlinear resonators external electric and magnetic fields can be used to tune the mode detuning via Zeeman and Rashba-Dresselhaus splitting for instance. The response is defined by the eigenvalue splitting $\Delta_{\omega}(\epsilon,\alpha)$ from the exceptional point eigenvalue $\mu(\epsilon=0,\alpha)$ in dependence of the perturbation strength:
@@ V Topology of Exceptional Rings
In the literature, ERs are intensively studied for their non-trivial band topologies. To validate this feature for our linear ER, we calculate the $1^{\mathrm{st}}$ Chern number of the ER on the discretized $k$-space grid using the $U(1)$ link variable to calculate the Berry curvature $\Omega_{n}$ [90] for each lattice site k
@@
The Chern number is evaluated on closed circular surfaces, one located inside the ER as well as one enclosing the ER. In the latter case, consistent with findings in other physical systems featuring ERs [91, 92], our ER exhibits a Chern number of $C=\pm 1$, depending on the eigenvalue surface under investigation, numerically accurate to the $16^{\mathrm{th}}$ decimal place, confirming a topologically nontrivial structure. In the nonlinear regime, we compute the Chern number on three types of closed circular surfaces: (i) inside the inner ring, (ii) enclosing only the inner ring, and (iii) enclosing both the inner and outer rings. Both the EAR and the EXR exhibit $1^{\mathrm{st}}$ Chern numbers of $C=\pm 1$. The surface inside each ring amounts to $C=0$. For weak nonlinearity, the proximity of the two rings leads to a significant overlap of their Berry curvature distributions. Consequently, integrating over the surface enclosing only the inner ring yields finite Chern number values on the order of $\approx 1e^{-1}$ for the outer ring, despite the integration surface lying entirely inside the ring.
@@
Figure 7: Complex-valued energy spectrum with exceptional rings for finite Zeeman splitting in nonlinear regime. (a,c) Real- and imaginary part of eigenvalues $\mu$. (b,d) Cross sections of $k$-isotropic complex-valued energy spectrum in (a,c). The main panels show the real and imaginary parts of the energy spectrum and central cuts that host the two concentric exceptional arc rings. The turquoise line in the inset in (b) shows the cross-section of the umbilic for finite $\alpha=0.1~{}\mathrm{meV\upmu m^{2}}$. The orange line displays the Zeeman splitting-induced detuning (increasing with increasing $\Delta_{\mathrm{B}}$). The displayed exceptional rings form at radii defined by the intersections with the two exceptional arcs. The coloring in (b) and (d) illustrates the mode polarization in form of the $S_{3}$ Stokes vector element for positive $\Delta_{\mathrm{B}}$.
@@
As also discussed in Ref. [91] the ER only exists in the eigenvalues of the matrix in Eq. (1) if the two modes are degenerate in energy ($\Delta=0$) over all k in the vicinity of the ER [Fig. 6(a)]. In microcavities, however, different physical effects (e.g., Zeeman and Rashba-Dresselhaus splitting) can induce different forms of, even k-dependent, detuning, which adds dimensions to the parameter space that can be accessed and explored. To show that the ER is not destroyed with finite detuning but instead can be recovered in a higher-dimensional parameter space, we extend the linear eigenvalue problem (1) by finite (k-dependent) detuning (with Zeeman and Rashba-Dresselhaus splittings as our specific examples of physical causes of this detuning). To this end $\Delta_{\pm}=\pm\Delta_{\mathrm{B}}\pm\Delta_{\mathrm{RB}}(k_{\parallel})$ is added to the first (+) and second (-) diagonal elements in Eq. (1). Here $\Delta_{\mathrm{B}}$ describes the magnetic field induced Zeeman splitting and $\Delta_{\mathrm{RB}}=2\alpha_{\mathrm{RD}}k_{\parallel}$ the Rashba-Dresselhaus splitting [93, 94, 95] with splitting strength $\alpha_{\mathrm{RD}}$. In specifically designed semiconductor microcavities $\alpha_{\mathrm{RD}}$ is adjustable by an external electric field acting on an optically anisotropic liquid crystal inside the cavity [20, 93, 94, 96]. Here we denote the Rashba-Dresselhaus splitting direction in $k$-space as $k_{\parallel}$ and the direction perpendicular to the splitting as $k_{\perp}$. We emphasize that the exact mathematical form and physical origin of detuning is not important for our results, those transfer to any form of detuning breaking $k$-isotropy in a similar fashion.
@@
By introducing the Zeeman shift $\Delta_{\mathrm{B}}$ to the circularly polarized modes via a magnetic field (which may be enabled with an optically active matter component inside the resonator), the two modes become non-degenerate and the ER vanishes [Fig. 6(b)]. As discussed, in certain structures and materials, an $E$-field induced Rashba-Dresselhaus splitting $\Delta_{\mathrm{RD}}$ along $k_{\mathrm{\parallel}}$ may be used to lift the degenercy of TE and TM modes in $k_{\mathrm{\parallel}}$-direction [95]. Thus, in the $k_{\mathrm{x}}$-$k_{\mathrm{y}}$ plane only the exceptional point pair along $k_{\mathrm{\perp}}=0$ remains visible [Fig. 6(c)]. The combination of these two effects lets us see the two exceptional points on the ring positioned where $k_{\mathrm{\perp}}\neq 0$ with $\Delta_{\mathrm{RD}}(k_{\parallel})-\Delta_{\mathrm{B}}=0$ [Fig. 6(d)]. With this approach, we can validate that the ER structure stays intact in higher-dimensional parameter space ($k_{\mathrm{x}},k_{\mathrm{y}},\Delta_{\mathrm{B}},\alpha_{\mathrm{RD}}$) even for finite detuning. In Fig. 6(e) the ER is shown in this 4D parameter space, the dots mark the location of exceptional points and thus the points for which the Zeeman shift compensates the Rashba-Dresselhaus splitting (note that the ERs form a continuous structure; for discretization reasonsn this is not fully resolved in the plot). Here, the tilt of the individual representations of the ER is induced by the respective Rashba-Dresselhaus splitting marked in color. While we are at present unable to calculate the Chern numbers for the tilted rings, we note that we expect each colored ring to possess a $1^{\mathrm{st}}$ Chern numbers of $C_{\mathrm{n}}=\pm 1$. Note that the radius of the ER projected onto the $k_{\mathrm{x}}$-$k_{\mathrm{y}}$ plane remains unchanged by these extensions. The detuning strengths chosen here and below are in agreement with possible Zeeman and Rashba-Dresselhaus splittings measured in microcavity systems with the perovskite semicondcutor $\mathrm{CsPbBr_{3}}$ as active medium [95, 97]. This observation shows that in higher dimensional parameter space the ER is part of an exceptional surface and the ring is resilient against detuning. While detuning values might not be fully tunable in experiments, this result illustrates that the linear ER can be (partially) recovered even in $k$-anisotropic samples. Note that this analysis extends to any similar form of detuning or anisotropy in this and other physical systems. Moreover, the tilt of the ER permits new ways for tracing trajectories in higher-order dimensional space, such as Solomon’s knots around the ER. This is promissing for both mode switching applications as well as investigation of topological characteristics, such as higher Chern numbers. In the following we will extend this investigation into the nonlinear regime, to show that the elliptic umbilic leads to even richer spectral topology and exceptional point structures in the nonlinear regime.

### v2

@@ III Nonlinearity-induced third-order exceptional ring
We begin the discussion of the nonlinear eigenvalue problem for the case of zero detuning, $\Delta=0$, showing eigenvalues and further analysis in Figs. 3 and 4. As shown in Appendix B, we have two solutions in region A and C, and four solutions in region B. In the following figures, for clarity, we show the eigenvalues $\mu^{\prime}=\mu-\frac{\hbar^{2}\bm{k}^{2}}{2m}$, where the parabolic dispersion is substracted. For a better comparison with the results in Section II, we set $\{\Gamma_{+},\Gamma_{-}\}=\{-0.2,-0.4\}~\mathrm{ps^{-1}}$ and $\Delta_{\mathrm{LT}}=0.1~\mathrm{meV\upmu m^{2}}$ for the remainder of this work. Figure 3 shows the resulting real and imaginary parts of the complex-valued energy spectrum and central cuts for $\{\alpha,g_{\mathrm{c}},g_{\mathrm{x}}\}=\{0.125,0.25,0\}~\mathrm{meV\upmu m}^{2}$ [see Fig. 3(c-f)] in comparison to the linear results [see Fig. 3(a-b)]. In the nonlinear regime the complex-valued energy spectrum hosts two concentric rings marked by solid blue and violet dots in the central cuts in Fig. 3(c-d), and labled by EAR and EXR in Fig. 3(e-f), separating three radial segments (A-C). The inner segment $A$ hosts two circularly polarized solutions, the segment $B$ between the rings, marked in red, features four solutions (two circularly and two linearly polarized), and the outer segment $C$ shows two linearly polarized solutions. Note, that nonlinearity shifts the non-linearly polarized real parts of the solutions in segment $A$ and $B$ by $2\alpha$, while the linearly polarized solutions in segment $B$ and $C$ are shifted by $\alpha$. In Figure 4(a-b) these segments are sketched for an arbitrary nonlinearity $\alpha\neq 0$ to visualize the bifurcation behavior at the two rings. Analyzing this behavior, we conclude that in the nonlinear regime the linear ER splits into an inner ring of fold points (blue dot) along $|\bm{k}|^{2}=\nicefrac{{|\gamma|}}{{|\Delta_{\mathrm{LT}}|}}$ and an outer ring of cusp points (violet dot) along $\sqrt{\nicefrac{{(\alpha^{2}+\gamma^{2})}}{{\Delta_{\mathrm{LT}}^{2}}}}$, which grows with increasing Kerr-like nonlinearity $g_{\mathrm{c}}$ or cross-interaction strength $g_{\mathrm{x}}$. The fold and cusp points can be identified with exceptional points of $2^{\mathrm{nd}}$ and $3^{\mathrm{rd}}$-order, respectively. Hence, the outer ring is a $3^{\mathrm{rd}}$-order exceptional ring. The circles mark points where a solution crosses the boundary between two segments without tracing through an ER (e.g. the two green colored solutions form an ER at the blue solid dot, at the border of segment A and B, while the blue and red colored solutions cross the border of segment A and B at the blue circle without forming an ER). For $\alpha<0$ the cusp ring is redshifted in relation to the fold ring [this scenario is displayed in Figure 1(a)]. While previous work realized individual higher-order exceptional points in two-dimensional nonlinear parameter space and thus reduced the parameter tuning requirements [75, 74], the ring structure observed here reduces these parameter tuning requirements even further, making higher-order exceptional point structures in low dimensions more accessible.
[…]
@@
Figure 4: Elliptic umbilic structure of exceptional rings and exceptional points in nonlinear regime. (a) Sketch of the three sections A-C [as displayed in Fig. 3(c)] with two, four, and two solutions, respectively. (b) The bifurcation behavior on the edges of ring B characterizes the inner (outer) ring as a ring of fold (cusp) points. (c) Radii of fold rings (turquoise lines) and cusp-rings (violet dots) for degnerate modes (segments marked as above) and for finite mode detuning $\Delta$. The umbilic structure shows that fold-rings are part of exceptional arcs and cusp-rings are located at exceptional nexuses. (d) Cross section in (c) for variable nonlinearity $\alpha$. The depicted elliptic umbilic predicts exceptional ring radii and type in the full parameter space. The central organization point represents a point on the linear exceptional ring discussed in Section II.
@@
Figure 5: Enhanced perturbation response at nonlinear exceptional rings and points. In (a-b) a $3^{\mathrm{rd}}$-order exceptional point on the exceptional nexus ring [star in (d)] is traced for increasing $\alpha$. (a) The inverse signal-enhancement factor $(\mathrm{SEF})^{-1}$ decreases in the nonlinear regime. The coloring visualizes $\alpha$ in each curve (growing from black to yellow). (b) The maximum $\mathrm{SEF}_{\mathrm{max}}$ at the exceptional point as a function of $\alpha$. (c) Response $\Delta_{\omega}$ of the exceptional points on the exceptional arc ring (EAR) and exceptional nexus ring (EXR) on the elliptic umbilic along $\Delta=0$ [horizontal dashed line in (d) corresponds to $\epsilon=0$ in (c)] for a fixed nonlinearity $\alpha$. The response of the EAR significantly deviates from the square-root law of linear exceptional points. Here, small perturbations tune the system along the left exceptional arc of the umbilic [red line in (d)]. Larger perturbations tune it away from the arc due to its curvature [blue area in (d)] leading to growing response. (d) Colored cut through elliptic umbilic at finite nonlinearity $\alpha$. The coloring denotes the response function scaling factor $\eta$ when perturbing an exceptional point on the umbilic singularity structure. Detuning allows to realize different rings on the umbilic, which feature distinct perturbation responses. Perturbation from the lower to the upper arc (or vice versa) leads to characteristic square-root response $\eta_{1}$. Nexus perurbation leads to stronger response $\eta_{2}$. The perturbation $\epsilon$ is defined in relation to the respective point on the umbilic. Simulation parameters: $\{g_{\mathrm{c}},g_{\mathrm{x}}\}=\{2\alpha,0\}~\mathrm{meV\upmu m}^{2}$.
In previous work [38] it was mentioned that the linear ER does not exist in the solutions of the eigenvectors for finite mode detuning. Here we have shown that in the nonlinear regime new ERs arise even for finite mode detuning. We use this knowledge to analyze the perturbation response of nonlinear ERs in Section IV and their topological features under mode detuning in Section V. In Appendix D we detail that the elliptic umbilic structure applies to a wide range of non-Hermitian Hamiltonians when introducing nonlinearity. Therefore, the topolgoical features and perturbation response of nonlinear ERs on the umbilic we discuss below, are generic and can be applied to exceptional points in many other systems and constellations.
@@ IV Enhanced Perturbation Response at Nonlinear Exceptional Rings and Points
In Ref. [75, 92] it was proposed that nonlinear exceptional points can exhibit a complete basis of eigenstates. In laser systems, the Petermann factor can be investigated as a measure of noise enhancement. Based on this, it was proposed in Ref. [74] that nonlinear exceptional points can exhibit a divergent signal-to-noise ratio. More recently, however, it was shown that the conventional Petermann factor does not cover the interplay of noise and nonlinearity and that nonlinear exceptional points exhibit finite signal-to-noise ratio [92], with a generalized form of the Petermann factor at exceptional points derived in Ref. [93]. In the following, we discuss that nonlinear exceptional points possess adjustable and in some cases ultra-enhanced perturbation response. Our analysis covers the group of both real- and imaginary-valued smooth nonlinearities that lead to the elliptic umbilic exceptional point structure described above; here exemplarily evaluated for Kerr-type coherent nonlinearities. The elliptic umbilic structure significantly increases the complexity of an exceptional points neighborhood permiting sensor design for exceptional points at different positions on the singularity structure, which changes the response to perturbations not limited to $n^{\mathrm{th}}$-root laws.
@@
In Figure 5(b) the maximum of the signal-enhancement factor $\mathrm{SEF}_{\mathrm{max}}$ is plotted against $\alpha$. The maximum signal enhancement of $\mathrm{SEF}=1.9\times 10^{9}$ is found for a nonlinearity strength $\nicefrac{{(g_{\mathrm{c}}-g_{\mathrm{x}})}}{{2}}=68~\mathrm{\upmu eV\upmu m^{2}}$. Here, the signal enhancement scales with the perturbation as $\mathrm{SEF}\propto|\epsilon|^{-1.29}$. In contrast to an isolated linear exceptional point or ER, the elliptic umbilic allows for different exceptional points to be perturbed depending on the initial detuning $\Delta$. Since the perturbation $\epsilon$ is parallel to $\Delta$ this means that depending on which exceptional point we chose to perturb, the response depends on its position on the umbilic. In planar nonlinear resonators external electric and magnetic fields can be used to tune the mode detuning via Zeeman and Rashba-Dresselhaus splitting for instance. The response is defined by the eigenvalue splitting $\Delta_{\omega}(\epsilon,\alpha)$ from the exceptional point eigenvalue $\mu(\epsilon=0,\alpha)$ in dependence of the perturbation strength:
@@ V Topology of Exceptional Rings
In the literature, ERs are intensively studied for their non-trivial band topologies [94]. In the following, we discuss several related aspects for our system. We have evaluated the vorticity for bands $n=u$ (upper band) and $n=l$ (lower band) in the case without nonlinearity, defined in Ref. [95]
$\nu_{ul}=-\frac{1}{2\pi}\oint_{\cal{L}}\nabla_{k}{\rm arg}(\mu^{\prime}_{u}(\bm{k})-\mu^{\prime}_{l}(\bm{k}))\cdot d\bm{k}$ (7)
@@
[…]
In Appendix E we provide more details regarding the topological properties of the ER in the $k_{x}-k_{y}$ plane, i.e. using $\Delta=0$. We find that the closed-loop Berry phase and the integral over the $z$-componennt of the Berry curvature depend on the choice of eigenvectors (left or right) used to calculate the Berry connection. Next, we calculate an integral over the Berry curvature that we call a quasi-Chern integral
$C^{ab}_{\mathrm{n}}=\frac{1}{2\pi}\int\Omega_{n}^{ab}(\bm{k})d\textbf{S}\,.$ (8)
If we use only right eigenvectors ($ab=RR$), we find that the integral over the $z$-componennt of the Berry curvature exists and, without nonlinearity, is equal to $C^{RR}_{\mathrm{n}}=\pm 1$ if the integration area includes the ER. Note that for combinations of left and right eigenvectors (i.e. $ab=LR$ or $ab=RL$), self-orthognality leads to divergence of quasi-Chern integral at the ER. A detailed discussion of the scaling of the integral value is given in Appendix E.
With the objective of analyzing the nonlinear solutions, the analytical results for the quasi-Chern integral are first validated numerically. For this purpose, we discretize the $k$-space grid and use the $U(1)$ link variable to calculate the Berry curvature for each lattice site $\bm{k}$ following Ref. [97]:
$C^{ab}_{\mathrm{n}}=\frac{1}{2\pi}\sum_{k}\Omega_{n}^{ab}(\bm{k})\delta k^{2}.$ (9)
@@
In the nonlinear regime, we compute the quasi-Chern integral on three types of round disks: (i) inside the inner ring, (ii) enclosing only the inner ring, and (iii) enclosing both the inner and outer rings. To calculate the quasi-Chern integral of the EAR and the EXR the integral is evaluated over the eigenvalue surfaces forming the respective ring. Here, both rings exhibit quasi-Chern integral of $C^{RR}_{\mathrm{n}}=\pm 1$. For weak nonlinearity, the proximity of the two rings leads to a significant overlap of their Berry curvature distributions. Consequently, integrating over the disk enclosing only the inner ring yields finite quasi-Chern integral values on the order of $\approx 10^{-1}$ for the outer ring, despite the integration disk lying entirely inside the ring.
@@
Figure 7: Complex-valued energy spectrum with exceptional rings for finite Zeeman splitting in nonlinear regime. (a,c) Real- and imaginary part of eigenvalues $\mu$. (b,d) Cross sections of $k$-isotropic complex-valued energy spectrum in (a,c). The main panels show the real and imaginary parts of the energy spectrum and central cuts that host the two concentric exceptional arc rings. The turquoise line in the inset in (b) shows the cross-section of the umbilic for finite $\alpha=0.1~\mathrm{meV\upmu m^{2}}$. The orange line displays the Zeeman splitting-induced detuning (increasing with increasing $\Delta_{\mathrm{B}}$). The displayed exceptional rings form at radii defined by the intersections with the two exceptional arcs. The coloring in (b) and (d) illustrates the mode polarization in form of the $S_{3}$ Stokes vector element for positive $\Delta_{\mathrm{B}}$. Simulation parameters: $\{\alpha,g_{\mathrm{c}},g_{\mathrm{x}}\}=\{0.1,0.2,0\}~\mathrm{meV\upmu m}^{2}$
As also discussed in Ref. [94] the ER only exists in the eigenvalues of the matrix in Eq. (1) if the two modes are degenerate in energy ($\Delta=0$) over all $\bm{k}$ in the vicinity of the ER [Fig. 6(a)]. Here, the eigenvalue problem (1) exibits passive $\mathcal{PT}$-symmetry (the system is $\mathcal{PT}$-symmetric up to a shift in eigenvalues), without detuning resulting in codimension of one for our exceptional points (which therefore form a gapless ring) [98]. Introducing detuning breaks the $\mathcal{PT}$-symmetry and increases the exceptional points’ codimension to two [73, 99], resulting in isolated exceptional points as reported in Ref. [94]. In microcavities, however, different physical effects (e.g., Zeeman and Rashba-Dresselhaus splitting) can induce different forms of, even $k$-dependent, detuning, which adds dimensions to the parameter space that can be accessed and explored.
@@
To show that the ER is not destroyed with finite detuning but instead can be recovered in a higher-dimensional parameter space, we extend the linear eigenvalue problem (1) by finite ($k$-dependent) detuning (with Zeeman and Rashba-Dresselhaus splittings as our specific examples of physical causes of this detuning). To this end $\Delta_{\pm}=\pm\Delta_{\mathrm{B}}\pm\Delta_{\mathrm{RB}}(k_{\parallel})$ is added to the first (+) and second (-) diagonal elements in Eq. (1). Here $\Delta_{\mathrm{B}}$ describes the magnetic field induced Zeeman splitting and $\Delta_{\mathrm{RB}}=2\alpha_{\mathrm{RD}}k_{\parallel}$ the Rashba-Dresselhaus splitting [100, 101, 102] with splitting strength $\alpha_{\mathrm{RD}}$. In specifically designed semiconductor microcavities $\alpha_{\mathrm{RD}}$ is adjustable by an external electric field acting on an optically anisotropic liquid crystal inside the cavity [21, 100, 101, 103]. Here we denote the Rashba-Dresselhaus splitting direction in $k$-space as $k_{\parallel}$ and the direction perpendicular to the splitting as $k_{\perp}$. We emphasize that the exact mathematical form and physical origin of detuning is not important for our results, those transfer to any form of detuning breaking $k$-isotropy in a similar fashion.
@@
By introducing the Zeeman shift $\Delta_{\mathrm{B}}$ to the circularly polarized modes via a magnetic field (which may be enabled with an optically active matter component inside the resonator), the two modes become non-degenerate and the ER vanishes [Fig. 6(b)]. As discussed, in certain structures and materials, an electric field induced Rashba-Dresselhaus splitting $\Delta_{\mathrm{RD}}$ along $k_{\mathrm{\parallel}}$ may be used to lift the degenercy of TE and TM modes in $k_{\mathrm{\parallel}}$-direction [102]. Thus, in the $k_{\mathrm{x}}$-$k_{\mathrm{y}}$ plane only the exceptional point pair along $k_{\mathrm{\perp}}=0$ remains visible [Fig. 6(c)]. The combination of these two effects lets us see the two exceptional points on the ring positioned where $k_{\mathrm{\perp}}\neq 0$ with $\Delta_{\mathrm{RD}}(k_{\parallel})-\Delta_{\mathrm{B}}=0$ [Fig. 6(d)]. With this approach, we can validate that the ER structure stays intact in higher-dimensional parameter space ($k_{\mathrm{x}},k_{\mathrm{y}},\Delta_{\mathrm{B}},\alpha_{\mathrm{RD}}$) even for finite detuning. In Fig. 6(e) the ER is shown in this 4D parameter space, the dots mark the location of exceptional points and thus the points for which the Zeeman shift compensates the Rashba-Dresselhaus splitting (note that the ERs form a continuous structure; for discretization reasons this is not fully resolved in the plot). Here, the tilt of the individual representations of the ER is induced by the respective Rashba-Dresselhaus splitting marked in color. While we are at present unable to calculate the quasi-Chern intergrals for the tilted rings, we note that we expect each colored ring to possess a $1^{\mathrm{st}}$ quasi-Chern intergral of $C^{RR}=\pm 1$. Note that the radius of the ER projected onto the $k_{\mathrm{x}}$-$k_{\mathrm{y}}$ plane remains unchanged by these extensions. The detuning strengths chosen here and below are in agreement with possible Zeeman and Rashba-Dresselhaus splittings measured in microcavity systems with the perovskite semicondcutor $\mathrm{CsPbBr_{3}}$ as active medium [102, 104]. This observation shows that in higher dimensional parameter space the ER is part of an exceptional surface and the ring is resilient against detuning. Increasing both parameter space dimensionality and codimension of the exceptional points explains the occurance of the exceptional surface, as discussed in [99]. While detuning values might not be fully tunable in experiments, this result illustrates that the linear ER can be (partially) recovered even in $k$-anisotropic samples. Note that this analysis extends to any similar form of detuning or anisotropy in this and other physical systems. Moreover, the tilt of the ER permits new ways for tracing trajectories in higher-order dimensional space, such as Solomon’s knots around the ER. This is promising for both mode switching applications as well as investigation of topological characteristics. In the following we will extend this investigation into the nonlinear regime, to show that the elliptic umbilic leads to even richer spectral topology and exceptional point structures in the nonlinear regime.